Solve the system of equations using substitution.
y = 2x + 9
y = 3x + 10

Answer by Zucchini(70)

In substitution, you just substitute one variable for the other. So, since y=3x+10, we can substitute that in the other equation:
3x + 10 = 2x + 9
x = -9
The last part was just solving regularly. We have solved for x, now we need to solve for y. So, we just plug in for x and solve.
3(-9)+ 10 = y
-21 + 10 = y
-11 = y
So, x is -9 and y is -11, so the ordered pair is (-9, -11).